# Height filtrations and base loci on flag bundles over a curve

### v1

@@ 1.4. Movable cones
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by $\langle\alpha^{\vee},\cdot\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle>0$ for any $w\in W/W_{P}$ with $\ell(w)\geq n-k+1$.
In particular, we find that $\mathcal{L}_{\lambda}$ is $k$-movable $\Longleftrightarrow$ the $k$-th minimum of Zhang of $h_{\mathcal{L}_{\lambda}}$ is positive.
@@ 1.5. Example: Grassmann bundles
We provide here the computation of essential minimum and big cone on Grassmann bundles. Let $k$ be an algebraically closed field and $C/k$ be a curve with function field $K=k(C)$. Let $E$ be a vector bundle of rank $n$ on $C$ with Harder-Narasimhan filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $B$ be the subgroup of $\operatorname{GL}_{n}$ of all upper triangular matrices, and let $Q$ be the parabolic subgroup containing $B$ of type $(\operatorname{rank}(E_{1}),\operatorname{rank}(E_{2}),\dots,\operatorname{rank}(E_{r}))$. Here we say a parabolic subgroup containing $B$ is of type $(a_{1},\dots,a_{r})$ if it consists of upper block triguluar matrices of shape
@@
Let $F(E)$ be the frame bundle of $E$, which is a $\operatorname{GL}_{n}$-principal bundle. Then the canonical reduction of $F(E)$ is $F_{Q}$ to $Q$ where $F_{Q}$ is the $Q$-bundle of frames respecting the filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $\mu_{i}$ ($i=1,\dots,n$) be the $i$-th number in the sequence
@@
Let $\operatorname{Gr}_{r}(E)=F(E)/P$ be the Grassmann bundle of $r$-dimension quotients, where $P$ is the parabolic subgroup containing $B$ of type $(n-r,r)$. Identify the Weyl group of $G$ with $S_{n}$ and of $P$ with $S_{n-r}\times S_{r}$ as usual. Then $W/W_{P}=S_{n}\big{/}(S_{n-r}\times S_{r})$ can be identified with the multi-index set $\big{\{}(i_{1},\dots,i_{r}):1\leq i_{1}<i_{2}<\cdots<i_{r}\leq n\big{\}}$ via
@@
Let $h_{\mathcal{O}(1)}$ be the height function on $\operatorname{Gr}_{r}(E_{K})$ induced by the model $(\operatorname{Gr}_{r}(E),\mathcal{O}(1))$. By Theorem A, the successive minima are $\langle{\mathrm{deg}}(F_{Q}),I\det_{2}\rangle=\sum_{i\in I}\mu_{i}$. (Theorem A says the successive minima are $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W_{Q}\backslash W/W_{P}$, but these numbers are really just $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W/W_{P}$.) In particular, the longest element $I_{0}=(1,2,\dots,r)$ gives the essential minimum
$\zeta_{\text{ess}}=\langle{\mathrm{deg}}(F_{Q}),I_{0}\det\nolimits_{2}\rangle=\mu_{1}+\cdots+\mu_{r}.$
The only element in $\Delta\backslash\Delta_{P}$ is $\alpha=\lambda_{n-r}-\lambda_{n-r+1}$ and the function $\langle\alpha^{\vee},\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\langle\alpha^{\vee},\det_{2}\rangle<0$ and $f\longmapsto 0$. Let $I_{0}=(1,2,\dots,r)\in W/W_{P}$ be the longest element. Then the function $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\mu_{1}+\cdots+\mu_{r}$ and $f\longmapsto 1$. By Theorem C, the big cone is given by $\langle\alpha^{\vee},\cdot\rangle<0$ and $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle>0$, and one sees that it is the cone given by extremal rays $f$ and $\mathcal{O}(1)-(\mu_{1}+\dots+\mu_{r})f$.
1.6. Notations and conventions
In the following, let $k$ be an algebraically closed field of characteristic zero and $C/k$ be a projective smooth curve with function field $K=k(C)$. Let $G/k$ be a connected reductive group and denote the identity component of its center by $Z(G)$. Fix a Borel subgroup $B\subseteq G$ and a maximal torus $T\subseteq B$. Let $W$ be the Weyl group and $\Delta$ be the set of simple roots with respect to $(G,B,T)$. For any $\alpha\in\Delta$, we denote by $\alpha^{\vee}$ the corresponding simple coroot.
We shall consider only parabolic subgroup $P\subseteq G$ containing $B$. For any parabolic subgroup $P$, let $W_{P}\subseteq W$ be the Weyl group $W(L_{P})$ of the Levi factor $L_{P}\subseteq P$ and $\Delta_{P}\subseteq\Delta$ be the simple roots of $L_{P}$. For parabolic subgroups $P,Q\subseteq G$ and any double coset $w\in W_{Q}\backslash W/W_{P}$, set $\ell(w)=\max_{\sigma\in W_{Q}wW_{P}}\min_{\tau\in\sigma W_{P}}\ell(\tau)$, which equals to $\operatorname{dim}QwP/P$. Here by abuse of notations, $\ell(\tau)$ is the usual length of the Weyl element $\tau$.
For any linear algebraic group $\Gamma/k$, let $X(\Gamma)$ denote the character group $\mathrm{Hom}(\Gamma,\mathbb{G}_{m})$. For any $\mathbb{Z}$-algebra $A$, denote by $X(\Gamma)_{A}$ the tensor product $X(\Gamma)\otimes_{\mathbb{Z}}A$. For any linear functional $f$ on $X(\Gamma)$, we shall usually denote $f(\alpha)$ by $\langle f,\alpha\rangle$ for $\alpha\in X(\Gamma)$. For any $\alpha\in X(\Gamma)$, denote by $k_{\alpha}$ the one-dimensional algebraic representation on the vector space $k$ with $\Gamma$ acting by $\alpha$.
1.7. Organization of the paper
Let $F$ be a principal $G$-bundle on $C$ and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly anti-dominant character. Then $F/P$ is a flag bundle over $C$ and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle. Let $X=(F/P)_{K}$ and $h_{\mathcal{L}_{\lambda}}$ be the height function on $X$ induced by $(F/P,\mathcal{L}_{\lambda})$.
In Chapter $2$, we compute the height filtration and successive minima of $h_{\mathcal{L}_{\lambda}}$. In Chapter $3$, we enhance Zhang’s inequality into an equality by convex geometry. In Chapter $4$, we compute the augmented base loci of $\mathcal{L}_{\lambda}-tf$ on $F/P$. As a corollary we get the movable cones $\operatorname{Mov}^{k}(F/P)$.
@@ 2. Height filtrations and successive minima
2.1. Vector bundles, slope stability and Harder-Narasimhan filtration
Let $E$ be a vector bundle on $C$. The degree of $E$ is ${\mathrm{deg}}(E)={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E)={\mathrm{deg}}(E)/\operatorname{rk}(E)$. It is called slope semistable if for every subbundle $F\subseteq E$, $\mu(F)\leq\mu(E)$. This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
In general, there exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
@@ 2.1. Vector bundles, slope stability and Harder-Narasimhan filtration
This filtration is called the Harder-Narasimhan filtration of $E$ and we denote $\mu_{i}(E)=\mu(E_{i}/E_{i-1})$. The first and the last term are of particular interests. The first one $\mu_{1}(E)=\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$. The last one $\mu_{r}(E)=\mu(E/E_{r-1})$ is called minimal slope of $E$ and is also denoted by $\mu_{\min}(E)$.
@@ 2.2. Principal bundles
For any linear algebraic group $\Gamma/k$, a principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
@@
Attached to any principal $\Gamma$-bundle $F$, one has the degree map
@@
Let $H$ be a closed subgroup of a linear algebraic group $\Gamma/k$. A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\varphi)$ where $F_{H}$ is a principal $H$-bundle and $\varphi:F_{H}\times_{H}\Gamma\cong F$ is an isomorphism. By the universal property of the quotient $F/H$, the assignment to any section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
Let $F$ be a principal $G$-bundle on $C$ for the connected reductive group $G$. Note that for any parabolic subgroup $P\subseteq G$ with Levi subgroup $L_{P}$, the natural inclusion
@@
and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [23, §2.1.3]. Let $F_{P}$ be a reduction of $F$ to $P$. By applying the slope map to ${\mathrm{deg}}(F_{P})$ we can define $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle$ for any $\lambda\in X(T)$.
2.3. Semistability and canonical reduction
We mainly follow [5].
Definition 2.1 (semistable).
A principal $G$-bundle $F$ on $C$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
Definition 2.2 (canonical reduction).
Let $F$ be a principal $G$-bundle on $C$. A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
@@ 2.3. Semistability and canonical reduction
Behrend in [3] proved canonical reduction exists uniquely. When $G=\operatorname{GL}_{n}$, the above definition of semistability recovers Mumford’s slope stability of vector bundles and the above deinition of canonical reduction recovers the Harder-Narasimhan filtration of vector bundles.
2.4. Height function induced by $(F/P,\mathcal{L}_{\lambda}$)
Let $F$ be a principal $G$-bundle with canonical reduction $F_{Q}$ to some parabolic subgroup $Q\subseteq G$. Let $P\subseteq G$ be a parabolic subgroup. Set $\mathcal{X}=F/P$ and $X=\mathcal{X}_{K}$.
A character $\lambda:P\longrightarrow\mathbb{G}_{m}$ is called strictly anti-dominant if the natural pairing $\langle\alpha^{\vee},\lambda\rangle<0$ for any $\alpha\in\Delta\backslash\Delta_{P}$. Let $\lambda:P\longrightarrow\mathbb{G}$ be a stricly anti-dominant character. Then the line bundle $M_{\lambda}=G\times_{P}k_{\lambda}$ on $G/P$ is ample. Therefore $\mathcal{L}_{\lambda}=F\times_{G}M_{\lambda}$ is a relatively ample line bundle on $\mathcal{X}=F\times_{G}G/P$ and induces a height function
$\quad\quad\quad h_{\mathcal{L}_{\lambda}}:X(\overline{K})$ $\longrightarrow\mathbb{Q},\quad x\longmapsto\frac{\mathcal{L}_{\lambda}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}$
where $\overline{\{x\}}$ is the Zariski closure of $x$ in $\mathcal{X}$.

### v2

@@ 1.3. Base loci and movable cones
The height filtration $Z_{t}$ in Theorem A measures the positivity of the height function $h_{\mathcal{L}_{\lambda}}$, and we see that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is just the Zariski closure of $Z_{t}$ in $\mathcal{X}$ for every $t\not=\zeta_{w}$. At critical points, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is upper continuous while $Z_{t}$ is lower continuous.
A class $\alpha\in N^{1}(F/P)_{\mathbb{R}}$ is called $k$-movable if $\mathrm{B}_{+}(\alpha)$ has codimension $\geq k$. As a direct corollary of Theorem D, we compute the movable cones on flag bundles. (Please see §4.3 for the notations.)
Theorem E (Theorem 4.2).
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ coincides with the cone defined by
  1. (1)
  2. (2)
Let $E$ be a vector bundle on a curve $C$. The ample cone of $\mathbb{P}(E)$ is computed in [21] and the big cone of $\mathbb{P}(E)$ is computed in [12]. Fulger and Lehmann computed $\operatorname{Mov}^{2}(\mathbb{P}(E))$ in [13]. In [4], the big cone of the Grassmann bundle of $E$ is computed and in [5], the ample cone of flag bundles of $E$ is computed. Our Theorem E is a simultaneous generalization of all these results.
Theorem F (Corollary 4.3).
Let $\mathcal{L}$ be a line bundle on $F/P$. Then $\mathcal{L}$ is $k$-movable if and only if $e_{k}(h_{\mathcal{L}})>0$. Here $e_{k}(h_{\mathcal{L}})$ is the $k$-th minimum of Zhang of the height function $h_{\mathcal{L}}$.
Take $k=1$, we recover [1, Theorem 1.1] (in this special case) that $\mathcal{L}_{\lambda}$ big $\Longleftrightarrow$ $\zeta_{\operatorname{ess}}(h_{\mathcal{L}_{\lambda}})>0$.
@@ 1.4. Example: Grassmann bundles
We provide here the computation of the essential minimum and the big cone on Grassmann bundles.
Let $k$ be an algebraically closed field and $C$ be a curve over $k$ with function field $K=k(C)$. Let $E$ be a vector bundle of rank $n$ on $C$ with Harder-Narasimhan filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $B$ be the subgroup of $\operatorname{GL}_{n}$ of all upper triangular matrices, and let $Q$ be the parabolic subgroup containing $B$ of type $(\operatorname{rank}(E_{1}),\operatorname{rank}(E_{2}),\dots,\operatorname{rank}(E_{r}))$. Here we say a parabolic subgroup containing $B$ is of type $(a_{1},\dots,a_{r})$ if it consists of upper block trianguluar matrices of the shape
@@
Let $F(E)$ be the frame bundle of $E$, which is a $\operatorname{GL}_{n}$-principal bundle. Then the canonical reduction of $F(E)$ to $Q$ is $F_{Q}$, the $Q$-bundle of frames respecting the filtration $0\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{r}$. Let $\mu_{i}$ ($1\leq i\leq n$) be the $i$-th number in the sequence
@@
Let $\operatorname{Gr}_{r}(E)=F(E)/P$ be the Grassmann bundle of $r$-dimension quotients, where $P$ is the parabolic subgroup containing $B$ of type $(n-r,r)$. Identify the Weyl group of $G$ with $S_{n}$ and of $P$ with $S_{n-r}\times S_{r}$ as usual and $W/W_{P}=S_{n}\big{/}(S_{n-r}\times S_{r})$ with the multi-index set $\big{\{}(i_{1},\dots,i_{r}):1\leq i_{1}<i_{2}<\cdots<i_{r}\leq n\big{\}}$ via
@@
Let $h_{\mathcal{O}(1)}$ be the height function on $\operatorname{Gr}_{r}(E_{K})$ induced by the model $(\operatorname{Gr}_{r}(E),\mathcal{O}(1))$. Let $I_{0}=(1,2,\dots,r)\in W/W_{P}$ be the longest Weyl element. By Theorem A, the essential minimum of $h_{\mathcal{O}(1)}$ is $\zeta_{\operatorname{ess}}=\langle{\mathrm{deg}}(F_{Q}),I_{0}\det_{2}\rangle=\mu_{1}+\cdots+\mu_{r}$.
The only element in $\Delta\backslash\Delta_{P}$ is $\alpha=\lambda_{n-r}-\lambda_{n-r+1}$ and the function $\langle\alpha^{\vee},\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\langle\alpha^{\vee},\det_{2}\rangle<0$ and $f\longmapsto 0$. $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle$ on $N^{1}(\operatorname{Gr}_{r}(E))$ is given by $\mathcal{O}(1)\longmapsto\mu_{1}+\cdots+\mu_{r}$ and $f\longmapsto 1$. By Theorem F, the big cone is given by $\langle\alpha^{\vee},\cdot\rangle<0$ and $\langle{\mathrm{deg}}(F_{Q}),I_{0}\cdot\rangle>0$. One sees that it is the cone given by extremal rays $f$ and $\mathcal{O}(1)-(\mu_{1}+\dots+\mu_{r})f$.
1.5. Acknowledgements
We thank Xinyi Yuan and Ye Tian for their consistent supports. We thank Jie Liu, Marc Besson, Hongsheng Hu, Mihai Fulger and Brian Lehmann for discussions and email correspondences.
The idea of comparing height filtrations and base loci was proposed by Huayi Chen in a private conversation. The idea of computing concave transform was inspired from discussions with François Ballaÿ. The idea of proving Proposition 3.4 by Weyl symmetry was inspired from discussions with Tao Gui. We wish to thank them especially.
@@ 2. Height filtrations and successive minima
Let $k$ be an algebraically closed field of characteristic zero and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$. Let $G$ be a connected reductive group over $k$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$. Let $\mathcal{X}=F/P$ and $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function.
${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. The main result in this section is the following theorem:
Theorem 2.1.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
Corollary 2.2.
The filtration $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ is finite, and is given by successively deleting Schubert cells $C_{w}$ with jumping points $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ for $w\in W_{Q}\backslash W/W_{P}$.
Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points $\zeta_{w}$ the successive minima. Note that our definition of successive minima are slightly different with Zhang [25]. Let $d=\dim G/P$, Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In the following, $e_{i}$ will be called Zhang’s successive minima to avoid ambiguity.
For $w\in W_{Q}\backslash W/W_{P}$, let
$\ell(w)=\max_{\sigma\in W_{Q}wW_{P}}\min_{\tau\in\sigma W_{P}}\ell(\tau)$
be the length function, which is nothing but the dimension of the cell $C_{w}$.
Corollary 2.3.
Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)\geq d-i+1\big{\}}$.
2.1. Vector bundles
Let $E$ be a vector bundle on $C$. The degree of $E$ is ${\mathrm{deg}}(E):={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E):={\mathrm{deg}}(E)/\operatorname{rk}(E)$. It is called slope semistable if for every subbundle $F$ of $E$, $\mu(F)\leq\mu(E)$. This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
There exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
  1. (1)
@@ 2.1. Vector bundles
This filtration is called the Harder-Narasimhan filtration of $E$. $\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$. $\mu(E/E_{r-1})$ is called minimal slope of $E$ and is also denoted by $\mu_{\min}(E)$.
@@ 2.2. Principal bundles
For any linear algebraic group $\Gamma/k$, let $X(\Gamma)=\mathrm{Hom}(\Gamma,\mathbb{G}_{m})$ denote the character group of $\Gamma$. For a cocharacter $f\in X(\Gamma)^{\vee}$ and a character $\lambda\in X(\Gamma)$, we shall denote the pairing by $\langle f,\lambda\rangle$. For any $\lambda\in X(\Gamma)$, denote by $k_{\lambda}$ the one-dimensional representation on the vector space $k$ with $\Gamma$ acting by $\lambda$.
A principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
@@
Attached to a principal $\Gamma$-bundle $F$, one has an associated cocharacter
@@
Let $H$ be a closed subgroup of a linear algebraic group $\Gamma$ over $k$. A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\phi)$ where $F_{H}$ is a principal $H$-bundle and $\phi:F_{H}\times_{H}\Gamma\simeq F$ is an isomorphism.
By the universal property of the quotient $F/H$, the assignment to a section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
2.3. Reductive groups, characters and cocharacters
Let $G$ be a connected reductive group over $k$. Fix a Borel subgroup $B\subseteq G$ and a maximal torus $T\subseteq B$. Let $W$ be the Weyl group and $\Delta$ be the set of simple roots with respect to $(G,B,T)$. For any $\alpha\in\Delta$, we denote by $\alpha^{\vee}$ the corresponding simple coroot.
We shall consider only parabolic subgroups containing $B$. For such a parabolic subgroup $P$, let $W_{P}\subseteq W$ be the Weyl group $W(L_{P})$ of the Levi factor $L_{P}\subseteq P$ and $\Delta_{P}\subseteq\Delta$ be the simple roots of $L_{P}$. Note that the natural inclusion
@@ 2.3. Reductive groups, characters and cocharacters
and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [23, §2.1.3]. In other words, a cocharacter on $X(P)_{\mathbb{Q}}$ can be extended canonically to $X(T)_{\mathbb{Q}}$.
2.4. The canonical reduction to a parabolic subgroup
Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$. Let $F_{P}$ be a reduction of $F$ to a parabolic subgoup $P$. Let ${\mathrm{deg}}(F_{P})\in X(T)_{\mathbb{Q}}^{\vee}$ be the induced (rational) cocharacter.
The $G$-bundle $F$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
Among all filtrations of a vector bundle, there is a canonical one (the Harder-Narasimhan filtration). Similarly, among all reduction to parabolic subgroups, there is a canonical one. A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
@@ 2.4. The canonical reduction to a parabolic subgroup
Behrend in [2] proved canonical reduction exists uniquely. Note that when $G=\operatorname{GL}_{n}$, the above definition of semistability recovers Mumford’s slope stability of vector bundles and the above deinition of canonical reduction recovers the Harder-Narasimhan filtration of vector bundles.
2.5. Arakelov geometry on flag varieties over function fields
Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $\mathcal{X}=F/P$ and $X=(F/P)_{K}$.
${X=(F/P)_{K},L\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P,\mathcal{L}_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$, $L=\mathcal{L}_{\lambda,K}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R},x\longmapsto\frac{\overline{\{x\}}\cdot\mathcal{L}_{\lambda}}{{\mathrm{deg}}(x)}$ be the induced height function. Here $\overline{\{x\}}$ is the Zariski closure of $x$ in $\mathcal{X}$.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Let $\zeta_{\operatorname{ess}}(X):=\inf\big{\{}t:Z_{t}=X\big{\}}$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X$.
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$.
$\mathcal{X}_{w}$ is a closed subscheme of $\mathcal{X}$, so $\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$ induces a height function $h_{\mathcal{L}_{\lambda}}:X_{w}(\overline{K})\longrightarrow\mathbb{R}$, which is nothing but the restriction of $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ to $X_{w}(\overline{K})$. Let $\zeta_{\operatorname{ess}}(X_{w})$ be the essential minimum of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$.
Theorem 2.1 follows from the following two propositions.
Proposition 2.4.
$h_{\mathcal{L}_{\lambda}}(x)\geq\langle{\mathrm{deg}}F_{Q},w\lambda\rangle$ for any $x\in C_{w}(\overline{K})$.
Proposition 2.5.
$\zeta_{\operatorname{ess}}(X_{w})=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
Theorem 2.1.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
Proof.
We write $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and $h=h_{\mathcal{L}_{\lambda}}$ to simplify notations. For a closed subvariety $Y$ of $X$, let $Z_{t}(Y)$ be the Zariski closure in $Y$ of $\big{\{}y\in Y(\overline{K}):h(y)<t\big{\}}$. By Proposition 2.4, $Z_{\zeta_{w}}(X_{w})\subseteq X_{w}\backslash C_{w}$, and we claim that $Z_{\zeta_{w}}(X_{w})=X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$.
In fact, suppose conversely that $Z_{\zeta_{w}}(X_{w})\subsetneq X_{w}\backslash C_{w}$. Then there exists $X_{w^{\prime}}\subsetneq X_{w}$ such that $Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. This forces $Z_{\zeta_{w}}(X_{w^{\prime}})\subseteq Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. But on $X_{w^{\prime}}$, Proposition 2.5 says the essential minimum is $\zeta_{w^{\prime}}<\zeta_{w}$, so $Z_{\zeta_{w}}(X_{w^{\prime}})=X_{w^{\prime}}$. Contradiction! Thus on $X_{w}$, the set $Z_{t}(X_{w})=X_{w}$ when $t>\zeta_{w}$ and $Z_{t}(X_{w})=X_{w}\backslash C_{w}$ when $t\leq\zeta_{w}$.
2.6. Proof of Proposition 2.4